My daughter is stuck on this equation. x+3x+(x-10)=180. I have the answer which is 38, but we have no idea how to get to this point. We tried using the reverse with the distributive property and grouping like terms and we are just totally lost. thanks for the help
Answer by SHUgrad05(58)

x+3x+(x-10)=180
Rewrite:
x+3x+x-10=180
Add 10 to both sides:
x+3x+x=190
Simplify:
5x=190
Divide:
x=38